on a map the scale is 1/2 inches =25 miles and I k=need to find the actual distance. Can you help me find the actual distance for 8 and 3/4inches?

Answer by checkley79(3341)

.5/8.75=25/X CROSS MULTIPLY.
.5X=8.75*25
.5X=218.75
X=218.75/.5
X=437.5 MILES.